Refer to Narrative 11-1. Calculate the compound interest on an investment of $45,000 at 6% interest, compounded quarterly, for 3 years.
Answer by Theo(13342)

you can use the formula f = p * (1+r) ^ n to solve this.
p is the present value = 45000
r = .06/4 = .015 per quarter
n = 3*4 = 12 quarters
formula becomes f = 45000 * 1.015 ^ 12 = 53802.82, rounded to the nearest penny.
the interest earned is that minus 45000 = 8802.82, rounded to the nearest penny.